### v1

# Geometric quantum discord of an arbitrary two-qudit state: the exact value and general upper bounds

Abstract
The geometric quantum discord of a two-qudit state has been studied in many papers, however, its exact analytical value in the explicit form is known only for a general two-qubit state, a general qubit-qudit state and some special families of two-qudit states. Based on the general Bloch vectors formalism [J. Phys. A: Math. Theor. 54 195301 (2021)], we find the explicit exact analytical value of the geometric quantum discord for a general two-qudit state of an arbitrary dimension via the parameters of its correlation matrix and the Bloch vectors of its reduced states. This new general analytical result indicates that the lower bound on the geometric quantum discord found in [Phys. Rev. A. 85, 204102 (2012)] is attained on each two-qudit state and also, includes all the known exact results on the geometric discord only as particular cases. Moreover, it allows us to find for an arbitrary two-qudit state, pure or mixed, the new general upper bounds on its geometric quantum discord, expressed via the Hilbert space characteristics of this state and in case of a pure two-qudit state – in terms of its concurrence.
1 Introduction
As proved by J. Bell theoretically [1] and later experimentally by A. Aspect et al. [2], some experiments exhibit correlations between outcomes observed by different parties under joint measurements that do not agree with the classical probability model, but are describable within the quantum formalism. However, not each multipartite quantum system in all experimental scenarios gives rise to nonclassical correlations – it is a particular property of the specific quantum systems, experimental arrangements, performed quantum measurements, etc. Nonclassicality of quantum correlations is the main resource in a variety of quantum information processing tasks.
Bell nonlocality and entanglement of a multipartite quantum state are the most studied of these quantum resources and are used [3] in many quantum information protocols and algorithms. For these concepts and the quantitative and qualitative relations between them, see [4, 5, 6, 7] and references therein.
Nevertheless, there are quantum states that exhibit nonclassical correlations even with small amounts of entanglement, and this led to the concept of the quantum discord [8] of a quantum state that quantifies a possible change of a quantum state under a measurement with the optimization over all possible measurements which could be performed on that state. This makes this concept conceptually rich but at the same time challenging, since it is very hard to compute it even for two qubit systems [9].
Due to the complexity of the computation of quantum discord, there are related concepts that have been proposed to quantify those quantum resources in an approximate way, although losing some of the advantages of the original concepts. This led to the appearance of some other measures quantifying nonclassicality of quantum correlations and the measurement-induced nonlocality [10] and the geometric quantum discord [11] are among them.
The geometric quantum discord is a geometric measure of quantum correlations, exhibited by a bipartite quantum state, which is defined via the distance from this state to the set $\Omega$ of all states that have the vanishing quantum discord and are referred to as classical-quantum [11] or quantum-classical [^1] In this paper, we refer to the right geometric discord instead of left discord as in [11], then the cited results will always be presented in the case of right geometric quantum discord.. This measure of quantum correlations is used in quantum phase estimation, quantum teleportation and remote state preparation protocols, see [12] and references therein.
In the present article, for the definition of the geometric quantum discord we consider the Hilbert-Schmidt distance between states. In this case, the geometric quantum discord $\mathcal{D}_{g}(\rho)$ of a two-qudit state $\rho$ on a complex Hilbert space $\mathcal{H}_{d_{1}}\otimes\mathcal{H}_{d_{2}}$ is defined by the relation
$\mathcal{D}_{g}(\rho):=\min_{\chi\in\Omega}||\rho-\chi||_{2}^{2},$ (1)
where $||\cdot||_{2}$ denotes the Hilbert-Schmidt norm. Possible other definitions [13, 14, 15] of the geometric quantum discord are based on a different than in (1) distance between quantum states.
Though the optimization problem involved into the computation of the geometric quantum discord is much simpler than that for the quantum discord, the geometric quantum discord of a bipartite state has been computed explicitly only in some particular cases, namely, for a general two-qubit state [11], a general qubit-qudit state [16] and some special families of two-qudit states [16, 17].
However, to our knowledge, the explicit exact analytical value of the geometric quantum discord for a general two-qudit state of an arbitrary dimension has not been reported in the literature – only its lower bounds, see in [16, 17, 18, 19, 20].
In the present paper, for an arbitrary two-qudit state $\rho$ on $\mathcal{H}_{d_{1}}\otimes\mathcal{H}_{d_{2}},$ pure or mixed, we find the explicit exact analytical value of its geometric quantum discord (1) via parameters of its correlation matrix and the Bloch vectors of its reduced states. This new result: (i) reduces in a two-qubit case to the well-known two-qubit expression of Dakić et al. [11]; (ii) proves that the lower bound found by Rana and Parashar in [16] is attained on each two-qudit state; (iii) allows us to find the new general upper bounds on the geometric quantum discord of an arbitrary two-qudit state in terms of its Hilbert space characteristics and in case of a pure two-qudit state – via its concurrence.
The new general results derived in the present article considerably extend the range of known results on properties of the geometric quantum discord of a two-qudit state, pure or mixed, of an arbitrary dimension.
The paper is organized as follows. In Section 2, we introduce the main issues of the general Bloch vectors formalism [22] for a finite-dimensional quantum system on which we build up the calculations in this paper. In Section 3, we find in the explicit analytical form the exact value of the geometric quantum discord for an arbitrary two-qudit state. In Section 4, this new result allows us to find new general upper and lower bounds on the geometric quantum discord in a general two-qudit case. In Section 5, we discuss the main results of this paper.
2 Preliminaries: general Bloch vectors formalism
In this section, we shortly recall the main issues of the general Bloch vectors mathematical formalism developed in [22] for the description of properties and behavior of a finite-dimensional quantum system.
Consider the vector space $\mathcal{L}_{d}$ of all linear operators $X$ on a complex Hilbert space $\mathcal{H}_{d}$ of a finite dimension $d\geq 2.$ Equipped with the scalar product $\langle X_{i},X_{j}\rangle_{\mathcal{L}_{d}}:=\mathrm{tr}\left[X_{i}^{{\dagger}}X_{j}\right],$ $\mathcal{L}_{d}$ is a Hilbert space of the dimension $d^{2},$ referred to as Hilbert-Schmidt. Denote by
$\mathfrak{B}_{\Upsilon_{d}}$ :=$:=\left\{\mathbb{I}_{d},\text{ }\Upsilon_{d}^{(k)}\in\mathcal{L}_{d},\text{ \ }k=1,...,(d^{2}-1)\right\},$ (2)
$\Upsilon_{d}^{(k)}$ $=\left(\Upsilon_{d}^{(k)}\right)^{\dagger}\neq 0,\ \ \ \mathrm{tr[}\Upsilon_{d}^{(k)}]=0,\ \ \ \mathrm{tr[}\Upsilon_{d}^{(k)}\Upsilon_{d}^{(m)}]=2\delta_{km},$
a basis of $\mathcal{L}_{d}$ consisting of the identity operator $\mathbb{I}_{d}$ on $\mathcal{H}_{d}$ and a tuple
$\Upsilon_{d}:=\left(\Upsilon_{d}^{(1)},...,\Upsilon_{d}^{(d^{2}-1)}\right)$ (3)
of traceless Hermitian operators mutually orthogonal in $\mathcal{L}_{d}.$ For $d\geq 3,$ some properties of a particular basis of this type, resulting in the generalized Gell-Mann representation, were considered in [23, 24, 25, 26, 27, 28, 29, 30, 31].
For an arbitrary qudit state $\rho$, the decomposition via a basis (2) constitutes the generalized Bloch representation [22]
$\rho_{d}$ $=\frac{\mathbb{I}_{d}}{d}\text{ }\mathbb{+}\text{ }\sqrt{\frac{d-1}{2d}}\left(r\cdot\Upsilon_{d}\right),\text{ \ \ \ }r\cdot\Upsilon_{d}:=\sum_{j=1}^{d^{2}-1}p_{j}\Upsilon_{d}^{(j)},$ (4)
 $\text{\ }r$ $=\sqrt{\frac{d}{2(d-1)}}\text{ }\mathrm{tr}\left[\rho_{d}\Upsilon_{d}\right]\in\mathbb{R}^{d^{2}-1},$ (5)
where $r\in\mathbb{R}^{d^{2}-1}$ is referred to as the Bloch vector of a qudit state $\rho_{d},$ and satisfies the relations
$\left\|r\right\|_{\mathbb{R}^{d^{2}-1}}^{2}=\frac{d}{d-1}\left(\mathrm{tr}\left[\rho_{d}^{2}\right]-\frac{1}{d}\right)\text{ \ \ }\Rightarrow\text{ \ }\left\|r\right\|_{\mathbb{R}^{d^{2}-1}}^{2}\leq 1\ .$ (6)
If a state $\rho_{d}$ is pure, then the norm of its Bloch vector $\left\|r\right\|_{\mathbb{R}^{d^{2}-1}}^{2}=1$. However, in contrast to a qubit case, for an arbitrary $d>2$, the converse is not true and not each unit vector $n\in\mathbb{R}^{d^{2}-1}$ corresponds via (4) to a pure state, see [22] for details.
As it is found in Lemma 1 of [31], the operator norm $\left\|\cdot\right\|_{0}$ of the operator $\left(r\cdot\Upsilon_{d}\right)$ satisfies the bounds
$\sqrt{\frac{2}{d}}\text{ \ }\leq\text{ \ }\frac{\left\|r\cdot\Upsilon_{d}\right\|_{0}}{\left\|r\right\|_{\mathbb{R}^{d^{2}-1}}}\text{ \ }\leq\text{ \ }\sqrt{\frac{2(d-1)}{d}}.$ (7)
For any two qudit states $\rho_{d},$ $\widetilde{\rho}_{d}$, the scalar product of their Bloch vectors satisfies the relation [22]
$r\cdot\widetilde{r}\geq-\frac{1}{d-1}\ ,$ (8)
where equality holds if $\mathrm{tr}[\rho_{d}\widetilde{\rho_{d}}]=0.$ Furthermore, for any orthonormal basis $\{|k\rangle\in\mathcal{H}_{d},$ $k=1,...,d\},$ the corresponding Bloch vectors $y_{k}\in\mathbb{R}^{d^{2}-1}$ of pure states $|k\rangle\langle k|$ of its elements obey the relations
$\sum_{k=1}^{d}y_{k}=0,\text{ \ }\left\|y_{k}\right\|=1\text{, \ \ \ }y_{k_{1}}\cdot y_{k_{2}}=-\frac{1}{d-1},\text{ \ \ \ }\forall k_{1}\neq k_{2}.$ (9)
For a two-qudit state $\rho_{d_{1}\times d_{2}}$ on $\mathcal{\ H}_{d_{1}}\otimes\mathcal{H}_{d_{2}},$ $d_{1},d_{2}\geq 2,$ the decomposition
$\rho_{d_{1}\times d_{2}}$ $=\frac{\mathbb{I}_{d_{1}}\otimes\mathbb{I}_{d_{2}}}{d_{1}d_{2}}\text{ }+\text{ }\sqrt{\frac{d_{1}-1}{2d_{1}}}\text{ }\left(r_{1}\cdot\Upsilon_{d_{1}}\right)\otimes\frac{\mathbb{I}_{d_{2}}}{d_{2}}\text{ }+\text{ }\sqrt{\frac{d_{2}-1}{2d_{2}}}\text{ }\frac{\mathbb{I}_{d_{1}}}{d_{1}}\otimes\left(r_{2}\cdot\Upsilon_{d_{2}}\right)$ (10)
[…]
via the basis of type (2) with the elements having the tensor product form
$\mathbb{I}_{d_{1}}\otimes\mathbb{I}_{d_{2}},\text{\ \ \ }\Upsilon_{d_{1}}^{(i)}\otimes\frac{\mathbb{I}_{d_{2}}}{\sqrt{d_{2}}},\text{ \ \ }\frac{\mathbb{\ I}_{d_{1}}}{\sqrt{d_{1}}}\otimes\Upsilon_{d_{2}}^{(j)},\text{ \ \ }\Upsilon_{d_{1}}^{(i)}\otimes\Upsilon_{d_{2}}^{(j)},$ (11)
$i$ $=1,...,(d_{1}^{2}-1),\text{ \ \ }j=1,...,(d_{2}^{2}-1),$
$\Upsilon_{d_{n}}^{(m)}$ $=\left(\Upsilon_{d_{n}}^{(m)}\right)^{\dagger}\neq 0,\ \ \ \mathrm{tr}\left[\Upsilon_{d_{n}}^{(m)}\right]=0,\ \ \ \mathrm{tr}\left[\Upsilon_{d_{n}}^{(m_{1})}\Upsilon_{d_{n}}^{(m_{2})}\right]=2\delta_{m_{1}m_{2}},\text{ \ \ }n=1,2,$
constitutes the generalized Pauli representation. In this representation [22],
$r_{1}$ $=\sqrt{\frac{d_{1}}{2(d_{1}-1)}}\mathrm{tr}\left[\rho_{d_{1}\times d_{2}}\left(\Upsilon_{d_{1}}\otimes\mathbb{I}_{d_{2}}\right)\right]\in\mathbb{R}^{d_{1}^{2}-1},\text{ \ \ }\left\|r_{1}\right\|_{\mathbb{R}^{d_{1}^{2}-1}}\leq 1,$ (12)
$r_{2}$ $=\sqrt{\frac{d_{2}}{2(d_{2}-1)}}\mathrm{tr}\left[\rho_{d_{1}\times d_{2}}\left(\mathbb{I}_{d_{1}}\otimes\Upsilon_{d_{2}}\right)\right]\in\mathbb{R}^{d_{2}^{2}-1},\text{ \ \ }\left\|r_{2}\right\|_{\mathbb{R}^{d_{2}^{2}-1}}\leq 1,$ (13)
[…]
$\mathrm{tr[}\rho_{j}^{2}]=\frac{1}{d_{j}}+\frac{d_{j}-1}{d_{j}}||r_{j}||_{\mathbb{R}^{d_{j}^{2}-1}}^{2},\text{ \ \ }j=1,2,$ (14)
while
$T_{\rho_{d_{1}\times d_{2}}}^{(ij)}=\mathrm{tr}\left[\rho_{d_{1}\times d_{2}}\left(\Upsilon_{d_{1}}^{(i)}\otimes\Upsilon_{d_{2}}^{(j)}\right)\right],\text{\ \ }i=1,...,d_{1}^{2}-1,\text{ \ \ }j=1,...,d_{2}^{2}-1,$ (15)
are the elements of the real-valued matrix $T_{\rho_{d_{1}\times d_{2}}}$ is called the correlation matrix of a two-qudit state $\rho_{d_{1}\times d_{2}}$.
In case of a pure two-qudit state $\rho_{d_{1}\times d_{2}}$ on $\mathcal{\ H}_{d_{1}}\otimes\mathcal{H}_{d_{2}},$ $d_{1},d_{2}\geq 2,$ by the Schmidt theorem $\rho_{1}^{2}]=\mathrm{tr}$[$\rho_{2}^{2}]$ and, in view of relation (14), this implies
$\frac{1}{d_{1}}+\frac{d_{1}-1}{d_{1}}||r_{1}||_{\mathbb{R}^{d_{1}^{2}-1}}^{2}=\frac{1}{d_{2}}+\frac{d_{2}-1}{d_{2}}||r_{{}_{2}}||_{\mathbb{R}^{d_{2}^{2}-1}}^{2}\ .$ (16)
From representation (10) and relation tr$[\rho_{d_{1}\times d_{2}}^{2}]\leq 1$ it follows
$\frac{d_{1}-1}{d_{1}d_{2}}\left\|r_{1}\right\|_{\mathbb{R}^{d_{1}^{2}-1}}^{2}+\frac{d_{2}-1}{d_{1}d_{2}}\left\|r_{2}\right\|_{\mathbb{R}^{d_{2}^{2}-1}}^{2}+\frac{1}{4}\sum_{i,j}\left(\ T_{\rho_{d_{1}\times d_{2}}}{}^{(ij)}\right)^{2}\leq\frac{d_{1}d_{2}-1}{d_{1}d_{2}},$ (17)
where equality holds in case of a pure state. Note that $\sum_{i,j}\left(T_{\rho_{d_{1}\times d_{2}}}^{(ij)}\right)^{2}$equals to the trace $\mathrm{tr[}T_{\rho_{d_{1}\times d_{2}}}^{\dagger}T_{\rho_{d_{1}\times d_{2}}}]$  of the positive operator $T_{\rho_{d_{1}\times d_{2}}}^{\dagger}T_{\rho_{d_{1}\times d_{2}}}$ on $\mathbb{R}^{d_{2}^{2}-1}.$
If $d_{1}=d_{2},$ then
$\left\|r_{1}\right\|_{\mathbb{R}^{d^{2}-1}}^{2}+\left\|r_{2}\right\|_{\mathbb{R}^{d^{2}-1}}^{2}+\frac{d^{2}}{4(d-1)}\sum_{i,j}\left(T_{\rho_{d\times d}}^{(ij)}\right)^{2}\leq d+1.$ (18)
The above inequalities constitute equalities if the two-qudit state $\rho_{d_{1}\times d_{2}}$ is pure.
Furthermore, for every two-qudit state $\rho_{d\times d}$ on $\mathcal{\ H}_{d}\otimes\mathcal{H}_{d},$ $d\geq 2,$ pure or mixed, from the bound (48) found in [31] and the above upper bound in (7) it follows that
$\left\|T_{\rho_{d\times d}}n\right\|_{\mathbb{R}^{d^{2}-1}}^{2}$ $\leq\sqrt{\frac{2}{d}}\sqrt{\frac{2(d-1)}{d}}\left\|T_{\rho_{d\times d}}n\right\|_{\mathbb{R}^{d^{2}-1}}\text{ \ }\Rightarrow\text{ \ }$
$\left\|T_{\rho_{d\times d}}n\right\|_{\mathbb{R}^{d^{2}-1}}$ $\leq 2\frac{\sqrt{d-1}}{d},$ (19)
for all $n\in\mathbb{R}^{d^{2}-1},\left\|n\right\|\leq\frac{1}{\sqrt{d-1}}.$ This implies that, for every two-qudit state, the spectral (operator) norm $\left\|T_{\rho_{d\times d}}\right\|_{0}$ of the correlation matrix $T_{\rho_{d\times d}}$ is upper bounded by
$\left\|T_{\rho_{d\times d}}\right\|_{0}$ $:=\sup_{\left\|n\right\|=1}\left\|T_{\rho_{d\times d}}n\right\|_{\mathbb{R}^{d^{2}-1}}=\sqrt{d-1}\sup_{\left\|n\right\|=1}\left\|T_{\rho_{d\times d}}(\frac{n}{\sqrt{d-1}})\right\|_{\mathbb{R}^{d^{2}-1}}$ (20)
$\leq\frac{2(d-1)}{d}.$
Recall that $\left\|T_{\rho_{d\times d}}\right\|_{0}^{2}$ is the maximal eigenvalue of the positive self-adjoint operator $T_{\rho_{d\times d}}^{\dagger}T_{\rho_{d\times d}}.$
For every separable two-qudit state
$\rho_{d\times d}^{(sep)}=\sum_{k}\beta_{k}\rho_{1}^{(k)}\otimes\rho_{2}^{(k)},\text{ \ \ }\beta_{k}\mathcal{\ >}0,\text{ \ \ }\sum_{k}\beta_{k}\mathcal{=}1,$ (21)
on $\mathcal{H}_{d}\otimes\mathcal{H}_{d},$ $d\geq 2,$ the correlation matrix $T_{\rho_{d\times d}^{(sep)}}$ and the Bloch vectors (12), (13) have the form
$T_{\rho_{d\times d}^{(sep)}}$ $=\frac{2(d-1)}{d}\sum_{k}\beta_{k}|r_{1}^{(k)}\rangle\langle r_{2}^{(k)}|,$ (22)
$r_{1}$ $=\sum_{k}\beta_{k}r_{1}^{(k)},\text{ \ \ }r_{2}=\sum_{k}\beta_{k}r_{2}^{(k)},$
where $r_{j}^{(k)}$ are the Bloch vectors of states $\rho_{j}^{(k)},$ $j=1,2,$ given by (5), and the operator norm of the correlation matrix is upper bounded by
$\left\|T_{\rho_{d\times d}^{(sep)}}\right\|_{0}$ $\leq\frac{2(d-1)}{d}\sum_{k}\beta_{k}\left\||r_{1}^{(k)}\rangle\langle r_{2}^{(k)}|\right\|_{0}$ (23)
$=\frac{2(d-1)}{d}\sum_{k}\beta_{k}\left\|r_{1}^{(k)}\right\|_{\mathbb{R}^{d^{2}-1}}\left\|r_{2}^{(k)}\right\|_{\mathbb{R}^{d^{2}-1}}$
$\leq\frac{2(d-1)}{d}\ .$
The concurrence $\mathrm{C}_{|\psi\rangle\langle\psi|}$ of a pure two-qudit state $\rho_{d_{1}\times d_{2}}=|\psi\rangle\langle\psi|$ on $\mathcal{H}_{d_{1}}\otimes\mathcal{H}_{d_{2}},$ $d_{1},d_{2}\geq 2,$ is defined by the relation
$\mathrm{C}_{|\psi\rangle\langle\psi|}=\sqrt{2\left(1-\mathrm{tr[}\rho_{j}^{2}]\right)}\ ,$ (24)
and, in view of Eqs. (14), (16), takes the form
$\mathrm{C}_{|\psi\rangle\langle\psi|}=\sqrt{2\frac{d_{j}-1}{d_{j}}\left(1-\left\|r_{j}\right\|_{\mathbb{R}^{d_{j}^{2}-1}}^{2}\right)},\text{ \ }\ j=1,2.$ (25)
If to introduce the concurrence $\widetilde{\mathrm{C}}_{|\psi\rangle\langle\psi|}$ normalized to the unity in case of a maximally entangled quantum state, as it was introduced in [22], then
$\mathrm{C}_{|\psi\rangle\langle\psi|}=\sqrt{2\frac{d_{k}-1}{d_{k}}}\widetilde{\mathrm{C}}_{|\psi\rangle\langle\psi|}\text{, \ \ \ \ \ }d_{k}\ =\min\{d_{1},d_{2}\},$ (26)
and
$\widetilde{\mathrm{C}}_{|\psi\rangle\langle\psi|}=\sqrt{1-\left\|r_{k}\right\|_{\mathbb{R}^{d_{k}^{2}-1}}^{2}}\ .$ (27)
In a two-qubit case, $\widetilde{\mathrm{C}}_{|\psi\rangle\langle\psi|}=\mathrm{C}_{|\psi\rangle\langle\psi|}.$ For a general state $\rho,$ pure or mixed, concurrence $\mathrm{C}_{\rho}$ is defined via the relation
$\mathrm{C}_{\rho}:=\inf_{\{\alpha_{i},\psi_{i}\}}\sum\alpha_{i}\mathrm{C}_{|\psi_{i}\rangle\langle\psi_{i}|},$ (28)
where $\rho=\sum\alpha_{i}|\psi_{i}\rangle\langle\psi_{i}|,$ $\sum\alpha_{i}=1\ ,$ $\alpha_{i}>0,$ is a possible convex decomposition of the state $\rho$ via pure states, see [6, 32] and references therein.
3 Geometric quantum discord
Let [^2] See footnote 1.
$\chi_{d_{1}\times d_{2}}^{(sep)}$ $=\sum_{k=1}^{d_{2}}\alpha_{k}\sigma_{k}\otimes|k\rangle\langle k|,\text{ \ \ }\alpha_{k}\geq 0,\text{ \ \ }\sum_{k}\alpha_{k}=1,$ (29)
$|k\rangle$ $\in\mathcal{H}_{d_{2}},\text{ }\ \text{\ }k=1,...,d_{2},\text{ \ \ }\langle k_{j_{1}}|k_{j_{2}}\rangle=\delta_{j_{1}j_{2}},\text{ \ \ }\sum_{k=1}^{d_{2}}\text{ }|k\rangle\langle k|\text{ }=\mathbb{I}_{d_{2}},$
be a quantum-classical state [33] on $\mathcal{H}_{d_{1}}\otimes\mathcal{H}_{d_{2}}.$
For short, we further omit the below indices at states indicating its dimensions at two sites.
In order to find the geometric quantum discord $\mathcal{D}_{g}(\rho):=\min_{\chi}\mathrm{tr}[(\rho-\chi)^{2}]$ of a state $\rho$, let us consider the decomposition of the difference between states $\rho$ and $\chi$ via their generalized Pauli representations (10). We have
$\rho-\chi$ $=\sqrt{\frac{d_{1}-1}{2d_{1}}}\left((r_{1}-\sum_{k}\alpha_{k}x_{k})\cdot\Upsilon_{d_{1}}\right)\otimes\frac{\mathbb{I}_{\mathcal{H}_{d_{2}}}}{d_{2}}$ (30)
$+\sqrt{\frac{d_{2}-1}{2d_{2}}}\text{ }\frac{\mathbb{I}_{\mathcal{H}_{d_{1}}}}{d_{1}}\otimes\left((r_{2}-\sum_{k}\alpha_{k}y_{k})\cdot\Upsilon_{d_{2}}\right)$
$+\frac{1}{4}\sum_{\begin{subarray}{c}i=1,...,d_{1},\\ j=1,...,d_{2}\end{subarray}}\left(T_{\rho}^{(ij)}-2\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}}\sum_{k}\alpha_{k}x_{k}^{(i)}y_{k}^{(j)}\right)\Upsilon_{d_{1}}^{(i)}\otimes\Upsilon_{d_{2}}^{(j)},$ (33)
where the Bloch vectors $r_{1}\in\mathbb{R}^{d_{1}^{2}-1}$ and $r_{2}\in\mathbb{R}^{d_{2}^{2}-1}$ and $T_{\rho}^{(ij)}$ are defined in (12), (13) and (15), respectively, and have norms $\left\|r_{1}\right\|_{\mathbb{R}^{d_{1}^{2}-1}}\leq 1,\left\|r_{2}\right\|_{\mathbb{R}^{d_{2}^{2}-1}}\leq 1,$ whereas
$x_{k}$ $=\sqrt{\frac{d_{1}}{2\left(d_{1}-1\right)}}\mathrm{tr}\left[\sigma_{k}\Upsilon_{d_{1}}\right]\in\mathbb{R}^{d_{1}^{2}-1},\text{ \ \ }k=1,...,d,\text{\ \ \ }\left\|x_{k}\right\|_{\mathbb{R}^{d_{1}^{2}-1}}\leq 1,$ (34)
$y_{k}$ $=\sqrt{\frac{d_{2}}{2\left(d_{2}-1\right)}}\langle k|\Upsilon_{d_{2}}|k\rangle\in\mathbb{R}^{d_{2}^{2}-1},\text{ \ \ }\left\|y_{k}\right\|_{\mathbb{R}^{d_{2}^{2}-1}}=1,$
are, correspondingly, the generalized Bloch vectors of states $\sigma_{k}$ and mutually orthogonal pure states $|k\rangle\langle k|,$ $\sum_{k=1}^{d_{2}}$ $|k\rangle\langle k|$ $=\mathbb{I}_{d_{2}}.$ By relation (9),
$\sum_{k=1}^{d_{2}}y_{k}=0,\text{ }\left\|y_{k}\right\|_{\mathbb{R}^{d_{2}^{2}-1}}=1,\text{ \ \ }y_{k}\cdot y_{k_{1}}=-\frac{1}{d_{2}-1}$ (35)
Eq. (30) implies
$\mathrm{tr}[\left(\rho-\chi\right)^{2}]$ $=\frac{d_{1}-1}{d_{1}d_{2}}\left\|r_{1}-\sum_{k}\alpha_{k}x_{k}\right\|_{\mathbb{R}^{d_{1}^{2}\text{ }-\text{ }1}}^{2}$ (36)
$+\frac{d_{2}-1}{d_{1}d_{2}}\left\|r_{2}-\sum_{k}\alpha_{k}y_{k}\right\|_{\mathbb{R}^{d_{2}^{2}\text{ }-\text{ }1}}$
$+\frac{1}{4}\sum_{i,j}\left(T_{\rho}^{(ij)}-2\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}}\sum_{k}\alpha_{k}x_{k}^{(i)}y_{k}^{(j)}\right)^{2}$
and, under conditions (35), relation (36) reduces to
$\mathrm{tr}[(\rho-\chi)^{2}]$ $=\frac{d_{2}-1}{d_{1}d_{2}}\left\|r_{2}\right\|_{\mathbb{R}^{d^{2}-1}}^{2}+\frac{1}{4}\mathrm{tr}[T_{\rho}^{t}T_{\rho}]$ (37)
$+\frac{d_{1}-1}{d_{1}}\sum_{k=1}^{d_{2}}\left\|\alpha_{k}x_{k}-\frac{r_{1}}{d_{2}}-\frac{1}{2}\sqrt{\frac{d_{1}\left(d_{2}-1\right)}{d_{2}\left(d_{1}-1\right)}}T_{\rho}y_{k}\right\|_{\mathbb{R}^{d^{2}-1}}^{2}$
$+\frac{1}{d_{1}}\sum_{k=1}^{d_{2}}\left[\alpha_{k}-\frac{1}{d_{2}}-\frac{d_{2}-1}{d_{2}}(r_{2}\cdot y_{k})\right]^{2}$
$-\frac{(d_{2}-1)^{2}}{d_{2}^{2}d_{1}}\sum_{k=1}^{d_{2}}(r_{2}\cdot y_{k})^{2}-\frac{d_{2}-1}{4d_{2}}\sum_{k=1}^{d_{2}}\left\|T_{\rho}y_{k}\right\|_{\mathbb{R}^{d_{2}^{2}-1}}^{2}\ .$
From relation (37) it follows that, for a fixed set $\Omega_{y},$ the minimum of $\mathrm{tr}[(\rho-\chi)^{2}]$ over $x_{k}$ and $\alpha_{k}$ is attained at
$\alpha_{k}x_{k}$ $=\frac{r_{1}}{d_{2}}-\frac{1}{2}\sqrt{\frac{d_{1}\left(d_{2}-1\right)}{d_{2}\left(d_{1}-1\right)}}T_{\rho}y_{k}$ (38)
$\alpha_{k}$ $=\frac{1}{d_{2}}+\frac{d_{2}-1}{d_{2}}(r_{2}\cdot y_{k})\Rightarrow\sum_{k=1}^{d_{2}}\alpha_{k}=1,$
such that
$\left\|\frac{r_{1}}{d_{2}}-\frac{1}{2}\sqrt{\frac{d_{1}\left(d_{2}-1\right)}{d_{2}\left(d_{1}-1\right)}}T_{\rho}y_{k}\right\|_{\mathbb{R}^{d_{1}^{2}-1}}\leq 1.$ (39)
Taking this into account in relations (37), we come to the following statement.
Proposition 1
For every two-qudit state $\rho$ on $\mathcal{H}_{d_{1}}\otimes\mathcal{H}_{d_{2}}$, the geometric quantum discord $\mathcal{D}_{g}(\rho)=\min_{\chi}\mathrm{tr}[(\rho-\chi)^{2}]$ is given by
$\mathcal{D}_{g}(\rho)$ $=\frac{d_{2}-1}{d_{1}d_{2}}\left\|r_{2}\right\|_{\mathbb{R}^{d_{2}^{2}-1}}^{2}+\frac{1}{4}\mathrm{tr}[T_{\rho}^{\dagger}T_{\rho}]$ (40)
$-\max_{\Omega_{y}}\mathrm{tr}\left[\left(\frac{d_{2}-1}{d_{1}d_{2}}|r_{2}\rangle\langle r_{2}|\text{ }+\text{ }\frac{1}{4}T_{\rho}^{\dagger}T_{\rho}\right)\Pi_{\Omega_{y}}\right],$
Here: (i) $T_{\rho}$ is the correlation matrix (15) of a state $\rho;$ (ii) $r_{2}$ is the Bloch vector (13) of the reduced state $\rho_{2}=\mathrm{tr}_{\mathcal{H}_{1}}[\rho]$ on $\mathcal{H}_{d_{2}}$; (iii) $\Pi_{\Omega_{y}}$ is the positive self-adjoint operator defined by the relation
$\Pi_{\Omega_{y}}:=\frac{d_{2}-1}{d_{2}}\sum_{k=1}^{d_{2}}|y_{k}\rangle\langle y_{k}|,$ (41)
where
$\Omega_{y}=\left\{y_{k}\in\mathbb{R}^{d_{2}^{2}-1},\text{ }k=1,..,d_{2}\right\}$ (42)
is the set of linear dependent Bloch vectors in $\mathbb{R}^{d_{2}^{2}-1},$ satisfying the relations
$\sum_{k=1}^{d_{2}}y_{k}=0,\text{ \ }\left\|y_{k}\right\|=1,\text{ \ }y_{k_{1}}\cdot y_{k_{2}}=-\frac{1}{d_{2}-1},\text{ \ }\forall k_{1}\neq k_{2}.$ (43)
In (40), (41), notations $|r_{2}\rangle$ and $\langle r_{2}|$ mean, respectively, the column vector and the line vector, corresponding to the tuple $r_{2}=(r_{2}^{(1)},...,r_{2}^{(d_{2}^{2}-1)})\in\mathbb{R}^{d_{2}^{2}-1}.$
In order to find via expression (40), the explicit analytical value of $\mathcal{D}_{g}(\rho)$ for an arbitrary two-qudit state, we formulate and prove in Appendix A the following new general statement.
Proposition 2
Let $A$ be a positive Hermitian operator on $\mathbb{R}^{d^{2}-1}$ and
$\Omega_{y}=\left\{y_{k}\in\mathbb{R}^{d^{2}-1},k=1,..,d\right\}$ (44)
be a set of linear dependent vectors in $\mathbb{R}^{d^{2}-1},$ satisfying relations (43). Then, for any set $\Omega_{y},$ the positive operator
$\Pi_{\Omega_{y}}=\frac{d-1}{d}\sum_{k=1}^{d}|y_{k}\rangle\langle y_{k}|$ (45)
on $\mathbb{R}^{d^{2}-1}$ is an orthogonal projection of rank $(d-1)$ and
$\max_{\Omega_{y}}\mathrm{tr}[A\Pi_{\Omega_{y}}]=\zeta_{1}+\cdots+\zeta_{d-1},$ (46)
where $\zeta_{1}\geq\zeta_{2}\geq\cdots\geq\zeta_{d^{2}-1}\geq 0$ are the eigenvalues of $A$ listed in the decreasing order with the corresponding algebraic multiplicities.
Taking into account Propositions 1 and 2, we derive for an arbitrary two-qudit state the explicit exact analytical value of its geometric quantum discord $\mathcal{D}_{g}(\rho)$ in terms of its characteristics in the generalized Pauli representation (10).
Theorem 1
For an arbitrary two-qudit state $\rho$ on $\mathcal{H}_{d_{1}}\otimes\mathcal{H}_{d_{2}},$ $d_{1},d_{2}\geq 2,$ the geometric quantum discord equals to
$\mathcal{D}_{g}(\rho)=\mathrm{tr}[G(\rho)]-\sum_{n=1}^{d_{2}-1}\eta_{n}=\sum_{n=d}^{d_{2}^{2}-1}\eta_{n},$ (47)
where
$G(\rho)=\frac{d_{2}-1}{d_{1}d_{2}}|r_{2}\rangle\langle r_{2}|\text{ }+\text{ }\frac{1}{4}T_{\rho}^{\dagger}T_{\rho}$ (48)
is a positive Hermitian operator on $\mathbb{R}^{d_{2}^{2}-1}$ and $\eta_{1}\geq\eta_{2}\geq\ldots\geq\eta_{d_{2}^{2}-1}\geq 0$  are its eigenvalues listed in the decreasing order with the corresponding algebraic multiplicities.
Theorem 1 proves that the lower bound on the geometric quantum discord presented in [16] is attained on any two-qudit state.
Note that, in contrast to the derivation of the lower bound in [16] via the generalized Gell-Mann representation, our derivation of the exact value (47) is based on any operator basis (11) and the normalization coefficients in (47) are different from those argued in [16] and satisfy the general relations derived in [22] and presented in short in Section 2.
Specified for a two-qubit case $d_{1}=d_{2}=2,$ the exact general result (47) leads to the known formula for the two-qubit geometric quantum discord found in [11].
Furthermore, our general result in Section 3 of [7] on the singular values of the correlation matrix of an arbitrary pure two-qubit state, allow us to derive based on (47) the following expression for the geometric quantum discord of an arbitrary pure two-qubit state via its concurrence and negativity without the assumption on the specific form [^3] For details, see Remark 1 in [7]. of a pure two-qubit state which is argued in [20].
Corollary 1
For any pure two-qubit state $|\psi\rangle\langle\psi|$ on $\mathcal{H}_{d_{1}}\otimes\mathcal{H}_{d_{2}},$ $d_{1},d_{2}\geq 2,$
$\mathcal{D}_{g}(|\psi\rangle\langle\psi|)=\frac{1}{2}\mathrm{C}_{|\psi\rangle\langle\psi|}^{2}=2\mathcal{N}_{|\psi\rangle\langle\psi|}^{2},$ (49)
where $\mathrm{C}_{|\psi\rangle\langle\psi|}$ is the concurrence (24) of a pure two-qubit state $|\psi\rangle\langle\psi|,$ and $\mathcal{N}_{|\psi\rangle\langle\psi|}$ is its negativity.
Proof. As it is proved found in Theorem 2 of [7], for a pure two-qubit state, the eigenvalues of the positive operator $T_{\rho}^{\dagger}T_{\rho}$ are equal to $1,$ $\mathrm{C}_{\rho}^{2}$, $\mathrm{C}_{\rho}^{2},$ where, for short, we use notation $\rho=|\psi\rangle\langle\psi|,$ and if $\left\|r_{2}\right\|_{\mathbb{R}^{3}}^{2}=1-\mathrm{C}_{\rho}^{2}\neq 0,$ that is, the state $|\psi\rangle$ is not maximally entangled, the vector $r_{2}\in\mathbb{R}^{3}$ constitutes the eigenvector of matrix $T_{\rho}^{\dagger}T_{\rho}.$ Therefore, if $|\psi\rangle$ is not maximally entangled, then, in view of the spectral theorem, the positive operator $G(\rho)$ in (48) takes the form
$G(\rho)$ $=\frac{1}{4}|r_{2}\rangle\langle r_{2}|+\frac{1}{4}\frac{|r_{2}\rangle\langle r_{2}|}{\left\|r_{2}\right\|_{\mathbb{R}^{3}}^{2}}+\frac{1}{4}\mathrm{C}_{\rho}^{2}\left(|\mathrm{v}_{1}\rangle\langle\mathrm{v}_{1}|+|\mathrm{v}_{2}\rangle\langle\mathrm{v}_{2}|\right)$ (50)
$=\frac{1}{4}\left(\left\|r_{2}\right\|_{\mathbb{R}^{3}}^{2}+1\right)\frac{|r_{2}\rangle\langle r_{2}|}{\left\|r_{2}\right\|_{\mathbb{R}^{3}}^{2}}+\frac{1}{4}\mathrm{C}_{\rho}^{2}\left(|\mathrm{v}_{1}\rangle\langle\mathrm{v}_{1}|+|\mathrm{v}_{2}\rangle\langle\mathrm{v}_{2}|\right)\ ,$
where $|\mathrm{v}_{1}\rangle,$ $|\mathrm{v}_{2}\rangle$ are two mutually orthogonal eigenvectors of $T_{\rho}^{\dagger}T_{\rho}$ corresponding to the eigenvalue $\mathrm{C}_{\rho}^{2}$ with multiplicity $2$. Representation (50) implies the eigenvalues of $G(\rho)$ are equal to
$\eta_{1}=\frac{1+\left\|r_{2}\right\|_{\mathbb{R}^{3}}^{2}}{4}\text{, }\eta_{2,3}=\frac{\mathrm{C}_{\rho}^{2}}{4}.$ (51)
If state $|\psi\rangle$ is maximally entangled, then $r_{2}=0,$ $T_{\rho}^{\dagger}T_{\rho}=\mathbb{I}_{2}$ and $G(\rho)=\frac{1}{4}\mathbb{I}_{2}.$ Therefore, for any pure state $|\psi\rangle,$ by (47) we have
$\mathcal{D}_{g}(\rho)=\sum_{n=2}^{3}\eta_{n}=\frac{1}{2}\mathrm{C}_{\rho}^{2}\ .$ (52)
The value of $\mathcal{D}_{g}(\rho)$ in terms of negativity follows from the relation $\mathrm{C}_{\rho}=2\mathcal{N}_{\rho}$ which is true for any pure two-qubit state.
In view of the eigenvalues of the correlation matrix for a two-qudit GHZ state found in [31], the exact formula (47) implies the following new result.
Corollary 2
For the GHZ state $\rho_{ghz}$ on $\mathcal{H}_{d}\otimes\mathcal{H}_{d},$ $d\geq 2,$
$\mathcal{D}_{g}(\rho_{ghz})=\frac{d-1}{d}.$ (53)
For any maximally entangled state pure two-qudit $\rho_{me}$ on $\mathcal{H}_{d_{1}}\otimes\mathcal{H}_{d_{2}},$ the geometric quantum discord is given by
$\mathcal{D}_{g}(\rho_{me})=\frac{d-1}{d},\text{ \ \ \ }d=\min\{d_{1},d_{2}\}.$ (54)
Proof. Taking into account that, as it is found in [31], for the two-qudit GHZ state $\rho_{ghz}$ on $\mathcal{H}_{d}\otimes\mathcal{H}_{d},$ $d\geq 2,$ (i) all singular values of its correlation matrix are equal to $\frac{2}{d},$ the reduced state $\rho_{2}=\frac{\mathbb{I}_{d}}{d},$ so that the Bloch vector $r_{2}=0$, we have
$G(\rho_{ghz})=\frac{1}{4}T_{\rho_{ghz}}^{\dagger}T_{\rho_{ghz}},\text{ \ \ \ }\eta_{n}=\frac{\lambda_{n}}{4}=\frac{1}{d^{2}}$ (55)
Therefore, by (47)
$\mathcal{D}_{g}(\rho_{ghz})=\sum_{n=d}^{d^{2}-1}\eta_{n}=\frac{d(d-1)}{d^{2}}=\frac{d-1}{d}.$ (56)
For any maximally entangled state pure state $\rho_{me}$ on $\mathcal{H}_{d_{1}}\otimes\mathcal{H}_{d_{2}},$ in view of its Schmidt decomposition and expression (53) for the GHZ state on $\mathcal{H}_{d}\otimes\mathcal{H}_{d},$, the geometric quantum discord is given by expression (54).
4 Upper and lower bounds
In this Section, based on the new result of Theorem 1, we introduce the general upper and lower bounds valid for any two-qudit state, also, specify them for a pure two-qudit state and a separable two-qudit state.
Theorem 2
For an arbitrary two-qudit state $\rho,$ pure or mixed, on $\mathcal{H}_{d_{1}}\otimes\mathcal{H}_{d_{2}},$ $d_{1},d_{2}\geq 2,$ its geometric quantum discord admits the following new upper bounds
$\mathcal{D}_{g}(\rho)$ $\leq\frac{d_{2}-1}{d_{2}}\left[1-\frac{\left\|r_{2}\right\|^{2}}{d_{2}+1}\right]$ (57)
$\leq\frac{d_{2}-1}{d_{2}},$ (58)
where the upper bound (58) constitutes the geometric discord of the maximally entangled two-qudit state if $\min\{d_{1},d_{2}\}=d_{2}.$
Proof. Taking into account that
$\mathrm{tr}[G(\rho)]=\sum_{n=1}^{d_{2}^{2}-1}\eta_{n}=\frac{d_{2}-1}{d_{2}d_{1}}\left\|r_{2}\right\|_{\mathbb{R}^{d_{2}^{2}-1}}^{2}+\frac{1}{4}\mathrm{tr}[T_{\rho}^{\dagger}T_{\rho}].$ (59)
and relation (17), we have
$\sum_{n=1}^{d_{2}^{2}-1}\eta_{n}\leq\frac{d_{1}d_{2}-1}{d_{1}d_{2}}-\frac{d_{1}-1}{d_{1}d_{2}}\left\|r_{1}\right\|_{\mathbb{R}^{d_{1}^{2}-1}}^{2}.$ (60)
We further note that since $\sum_{n=1}^{d_{2}^{2}-1}\eta_{n}=\sum_{n=1}^{d_{2}-1}\eta_{n}+$ $\sum_{n=d_{2}}^{d_{2}^{2}-1}\eta_{n}$ and $\sum_{n=d_{2}}^{d_{2}^{2}-1}\eta_{n}\leq d_{2}\sum_{n=1}^{d_{2}-1}\eta_{n}$, relation (60) implies
$\sum_{n=d_{2}}^{d_{2}^{2}-1}\eta_{n}$ $\leq\frac{1}{d_{2}+1}\left\{\frac{d_{1}d_{2}-1}{d_{1}}-\frac{d_{1}-1}{d_{1}}\left\|r_{1}\right\|{}_{\mathbb{R}^{d_{1}^{2}-1}}^{2}\right\}$ (61)
$=\frac{d_{1}d_{2}-1}{d_{1}(d_{2}+1)}-\frac{1}{d_{2}+1}\cdot\frac{d_{1}-1}{d_{1}}\left\|r_{1}\right\|^{2}.$
By using in (61) equality (16), we derive
$\sum_{n=d_{2}}^{d_{2}^{2}-1}\eta_{n}$ $\leq\frac{d_{1}d_{2}-1}{d_{1}(d_{2}+1)}-\frac{1}{d_{2}+1}\left(\frac{1}{d_{2}}-\frac{1}{d_{1}}+\frac{d_{2}-1}{d_{2}}\left\|r_{2}\right\|^{2}\right)$ (62)
$=\frac{d_{2}-1}{d_{2}}\left[1-\frac{\left\|r_{2}\right\|^{2}}{d_{2}+1}\right].$
The latter also implies the upper bound (58) and proves the statement.
For an arbitrary pure two-qudit state, relations (27), (57) and Corollary 2 imply the following new upper bound.
Corollary 3
For every pure two-qudit state $|\psi\rangle\langle\psi|$ on $\mathcal{H}_{d}\otimes\mathcal{H}_{d},$ $d\geq 2,$ the geometric quantum discord admits the following new upper bounds
$\mathcal{D}_{g}(|\psi\rangle\langle\psi|)$ $\leq\frac{d-1}{d(d+1)}\left(d+\widetilde{\mathrm{C}}_{|\psi\rangle\langle\psi|}^{2}\right)$ (63)
$\leq\frac{d-1}{d},$ (64)
where $\widetilde{\mathrm{C}}_{|\psi\rangle\langle\psi|}\leq 1$ is the normalized concurrence (26) of a pure two-qudit state $|\psi\rangle\langle\psi|$.
By Corollary 2, the upper bound (63) is attained on a maximally entangled pure two-qudit state.
Furthermore, based on the exact relation (47) we find the following general upper and lower bounds on the geometric quantum discord of a two-qudit state.
Theorem 3
For a two-qudit state $\rho$ on $\mathcal{H}_{d_{1}}\otimes\mathcal{H}_{d_{2}},$ $d_{1},d_{2}\geq 2,$ the geometric quantum discord admits the following new bounds:
$\frac{1}{4}\mathrm{tr}[T_{\rho}^{t}T_{\rho}]-\frac{1}{4}\sum_{n=1}^{d_{2}-1}\lambda_{n}$ (65)
$\leq\mathcal{D}_{g}(\rho)$
$\leq\min\left\{\frac{1}{4}\mathrm{tr}[T_{\rho}^{\dagger}T_{\rho}];\text{ }\frac{1}{4}\mathrm{tr}[T_{\rho}^{\dagger}T_{\rho}]+\frac{d_{2}-1}{d_{2}d_{1}}\left\|r_{2}\right\|_{\mathbb{R}^{d_{2}^{2}-1}}^{2}-\frac{1}{4}\sum_{n=1}^{d_{2}-1}\lambda_{n}\right\},$
where $\lambda_{1}\geq\lambda_{2}\geq\cdots\lambda_{d^{2}-1}\geq 0$ are eigenvalues of the positive Hermitian operator $T_{\rho}^{\dagger}T_{\rho}$ on $\mathbb{R}^{d_{2}^{2}-1}.$
Proof. For the evaluation of the last term in (40), we note that $\max_{x}\left\{f_{1}(x)+f_{2}(x)\right\}\leq\max_{x}f_{1}(x)+\max_{x}f_{2}(x),$ and, if $f_{j}(x)\geq 0,$ $j=1,2,$ then $\max_{x}\left\{f_{1}(x)+f_{2}(x)\right\}\geq\max_{x}f_{j}(x),$ $j=1,2.$ These relations and Proposition 2 imply
$\max_{\Omega_{y}}\mathrm{tr}\left[\left(\frac{d_{2}-1}{d_{1}d_{2}}|r_{2}\rangle\langle r_{2}|+\frac{1}{4}T_{\rho}^{\dagger}T_{\rho}\right)\Pi_{\Omega_{y}}\right]$ (66)
$\leq\max_{\Omega_{y}}\mathrm{tr}\left[\frac{d_{2}-1}{d_{1}d_{2}}|r_{2}\rangle\langle r_{2}|\Pi_{\Omega_{y}}\right]+\max_{\Omega_{y}}\mathrm{tr}\left[\frac{1}{4}T_{\rho}^{\dagger}T_{\rho}\Pi_{\Omega_{y}}\right]$
$=\frac{d_{2}-1}{d_{1}d_{2}}\left\|r_{2}\right\|_{\mathbb{R}^{d_{2}^{2}-1}}^{2}+\frac{1}{4}\sum_{n=1}^{d_{2}-1}\lambda_{n},$
as well as
$\max_{\Omega_{y}}\mathrm{tr}\left[\left(\frac{d_{2}-1}{d_{1}d_{2}}|r_{2}\rangle\langle r_{2}|+\frac{1}{4}T_{\rho}^{\dagger}T_{\rho}\right)\Pi_{\Omega_{y}}\right]$ (67)
$\geq\max_{\Omega_{y}}\mathrm{tr}\left[\left(\frac{d_{2}-1}{d_{1}d_{2}}|r_{2}\rangle\langle r_{2}|\right)\Pi_{\Omega_{y}}\right]=\frac{d_{2}-1}{d_{1}d_{2}}\left\|r_{2}\right\|_{\mathbb{R}^{d_{2}^{2}-1}}^{2}$
and
$\max_{\Omega_{y}}\mathrm{tr}\left[\left(\frac{d_{2}-1}{d_{1}d_{2}}|r_{2}\rangle\langle r_{2}|+\frac{1}{4}T_{\rho}^{\dagger}T_{\rho}\right)\Pi_{\Omega_{y}}\right]$ (68)
$\geq\max_{\Omega_{y}}\mathrm{tr}\left[\frac{1}{4}T_{\rho}^{\dagger}T_{\rho}\Pi_{\Omega_{y}}\right]=\frac{1}{4}\sum_{n=1}^{d_{2}-1}\lambda_{n}\ .$
Relations (66), (67) and (68) prove the statement.
From relation (17) and the upper bound in (65) it follows that, for a two qudit state $\rho$ on $\mathcal{H}_{d_{1}}\otimes\mathcal{H}_{d_{2}},$ $d_{1},d_{2}\geq 2,$
$\mathcal{D}_{g}(\rho)\leq\min\{J_{1},J_{2}\},$ (69)
where
$J_{1}=\frac{d_{1}d_{2}-1}{d_{2}d_{1}}-\frac{d_{1}-1}{d_{2}d_{1}}\left\|r_{1}\right\|_{\mathbb{R}^{d_{1}^{2}-1}}^{2}-\frac{d_{2}-1}{d_{2}d_{1}}\left\|r_{2}\right\|_{\mathbb{R}^{d_{2}^{2}-1}}^{2}$ (70)
and
$J_{2}=\frac{d_{1}d_{2}-1}{d_{2}d_{1}}-\frac{d_{1}-1}{d_{2}d_{1}}\left\|r_{1}\right\|_{\mathbb{R}^{d_{1}^{2}-1}}^{2}-\frac{1}{4}\sum_{n=1}^{d_{2}-1}\lambda_{n}.$ (71)
The new upper bounds (57), (58) in Theorem 2 and the new upper bounds (65), (69)-(71) considerably improve the upper bound in Proposition 3.1 of [19] having in our notations the form $\frac{d_{1}d_{2}-1}{d_{2}d_{1}}$.
Consider also the upper bound on the geometric quantum discord in a general separable case.
Proposition 3
For every separable two-qudit state
$\rho_{sep}=\sum_{k}\beta_{k}\rho_{1}^{(k)}\otimes\rho_{2}^{(k)},\text{ \ \ }\beta_{k}\mathcal{\ \geq}0,\text{ \ \ }\sum_{k}\beta_{k}\mathcal{=}1,$ (72)
on $\mathcal{H}_{d}\otimes\mathcal{H}_{d},$ $d\geq 2,$
$\mathcal{D}_{g}(\rho_{sep})\leq\left(\frac{d-1}{d}\right)^{2}.$ (73)
If a separable two qubit state is pure, then its geometric discord $\mathcal{D}_{g}(\rho_{sep}^{(pure)})=0.$
Proof. In view of expression (22) for the correlation matrix $T_{\rho_{sep}}$ of separable state $\rho_{sep},$ we have in case $d_{1}=d_{2}:$
$\frac{1}{4}\mathrm{tr}[T_{\rho_{sep}}^{\dagger}T_{\rho_{sep}}]$ $=\left(\frac{d-1}{d}\right)^{2}\ \sum_{k,k_{1}}\beta_{k}\beta_{k_{1}}(r_{1}^{(k)}\cdot r_{1}^{(k_{1})})(r_{2}^{(k)}\cdot r_{2}^{(k_{1})})$ (74)
$\leq\left(\frac{d-1}{d}\right)^{2},$
where $r_{j}^{(k)}$ are the Bloch vectors of states $\rho_{j}^{(k)},$ $j=1,2,$ given by (5) with norms $||r_{1}^{(k)}||,$ $||r_{2}^{(k)}||$ $\leq 1.$ If a separable state is pure, then $||r_{1}||=$ $||r_{2}||$ $=1$ and the positive Hermitian operator (48) is given by
$G(\rho_{sep}^{(pure)})$ $=\frac{d-1}{d^{2}}|r_{2}\rangle\langle r_{2}|\text{ }+\text{ }\left(\frac{d-1}{d}\right)^{2}\ \left\|r_{1}\right\|^{2}|r_{2}\rangle\langle r_{2}|$ (75)
$=\left(\frac{d-1}{d^{2}}+\left(\frac{d-1}{d}\right)^{2}\right)\text{ }|r_{2}\rangle\langle r_{2}|=\frac{d-1}{d}|r_{2}\rangle\langle r_{2}|$
and has only one nonzero eigenvalue
$\eta_{1}=\frac{d-1}{d}=\mathrm{tr}\left[G(\rho_{sep}^{(pure)})\right]$ (76)
with multiplicity one. Therefore, by (47) $\mathcal{D}_{g}(\rho_{sep}^{(pure)})=\mathrm{tr}\left[G(\rho_{sep}^{(pure)})\right]-\eta_{1}=0.$ This proves all items of the statement.
From the upper bound (73) it follows that, any separable two-qubit state, the geometric quantum discord cannot exceed $1/4$
5 Conclusion
In the present article, for an arbitrary two-qudit state with any dimensions at two sites, we find (Theorem 1) in the explicit analytical form the exact value of its geometric quantum discord. This new rigorously proved general result indicates that the lower bound found in [16] is attained on each two-qudit state and includes only as particular cases the exact results for a general two-qubit state [11], a general qubit-qudit state [16] and some special families of two-qudit states [16, 17].
Based on this new general result (47) of Theorem 1, we:
show (Corollary 1) that, for a pure two-qubit state, the exact value of the geometric quantum discord is proportional to its squared concurrence;
derive (Corollary 2) the exact value (54) of the geometric quantum discord for a maximally entangled pure two-qudit state of any dimension;
find (Theorem 2) the new general upper bounds (57) and (58) on the geometric quantum discord of an arbitrary two-qudit quantum state of any dimension which are consistent with the exact value in Corollary 2 for the geometric quantum discord of a maximally entangled pure two-qudit state.
specify (Corollary 3) for an arbitrary pure two-qudit state the new general upper bound (57) on its geometric quantum discord via the concurrence of this pure state. This upper bound is consistent with the exact value for a maximally entangled pure two-qudit state found in Corollary 2;
derive (Theorem 3) for an arbitrary two-qudit state the new general upper and lower bounds on the geometric quantum discord expressed via the eigenvalues of its correlation matrix. These upper bounds are tighter than the ones in [19].
specify the new upper bound on the geometric quantum discord of an arbitrary separable two-qudit state of any dimension.
The new general results derived in the present article considerably extend the range of known results on properties of the geometric quantum discord of a two-qudit state, pure or mixed, of an arbitrary dimension.
Acknowledgment
The study was implemented in the framework of the Basic Research Program at the HSE University.
6 Appendix A
First, let us show that every positive operator $\Pi_{\Omega_{y}}$ on $\mathbb{R}^{d^{2}-1}$, defined by relation (45), constitutes an orthogonal projection. In view of relations (43), we have
$\Pi_{\Omega_{y}}^{2}$ $=\left(\frac{d-1}{d}\right)^{2}\sum_{k,l\in\{1,...,d\}}|y_{k}\rangle\langle y_{k}|y_{l}\rangle\langle y_{l}|\text{ }$ (A1)
$=\left(\frac{d-1}{d}\right)^{2}\sum_{k=1,..,d}|y_{k}\rangle\langle y_{k}|\text{ }-\frac{d-1}{d^{2}}\sum_{k\neq l\in\{1,...,d\}}|y_{k}\rangle\langle y_{l}|\text{ }$
$=\left(\frac{d-1}{d}\right)^{2}\sum_{k=1,..,d}|y_{k}\rangle\langle y_{k}|\text{ }-\frac{d-1}{d^{2}}\sum_{k,l\in\{1,...,d\}}|y_{k}\rangle\langle y_{l}|\text{ }+\frac{d-1}{d^{2}}\sum_{k=1,..,d}|y_{k}\rangle\langle y_{k}|$
$=\frac{d-1}{d}\sum_{k=1,..,d}|y_{k}\rangle\langle y_{k}|\text{ }=\Pi_{\Omega_{y}}.$
Since
$\mathrm{tr}[\Pi_{\Omega_{y}}]=d-1,$ (A2)
the rank of every projection $\Pi_{\Omega_{y}}$ in Proposition 2 is equal to $\left(d-1\right).$ Note that an orthogonal projection $\Pi_{\Omega_{y}}$ has eigenvalue $1$ of multiplicity $(d-1)$ and eigenvalue 00 of multiplicity $d(d-1).$ Let
Denote by $e_{n},$ $n=1,...,N,$ mutually orthogonal unit eigenvectors of a positive Hermitian operator $A,$ corresponding to its eigenvalues $\zeta_{1}\geq\zeta_{2}\geq\cdots\geq\zeta_{N}\geq 0,$ listed in the decreasing order with the corresponding algebraic multiplicities,.and by $g_{m}\in\mathcal{H}_{N},$ $k=1,...,N,$ mutually orthogonal eigenvectors of projection $\Pi_{\Omega_{y}},$ where the first $(d-1)$ eigenvectors correspond to eigenvalue $1$. By the spectral theorem,
$A=\sum_{n=1}^{N}\zeta_{n}|e_{n}\rangle\langle e_{n}|,\text{ \ \ \ }\Pi_{\Omega_{y}}=\sum_{k=1}^{d-1}|g_{k}\rangle\langle g_{k}|.$ (A3)
By the von Neumann inequality [34] for the trace $\mathrm{tr}[AB]$ of the product of two positive operators, we have
$\mathrm{tr}[A\Pi_{\Omega_{y}}]\leq\sum_{k=1}^{d-1}\zeta_{k}.$ (A4)
Therefore, in order to prove relation (46) of Proposition 2, we have to present some projection $\Pi_{\widetilde{\Omega}_{y}}$ on which this upper bound is attained.
To a given projection $\Pi_{\Omega_{y}}=\frac{d-1}{d}\sum_{k=1}^{d}|y_{k}\rangle\langle y_{k}|,$ having the spectral decomposition (A3) and constructed via vectors in a set $\Omega_{y}=\left\{y_{k}\in\mathbb{R}^{d^{2}-1},\text{ }k=1,..,d\right\}$ satisfying relations (43), namely:
$\sum_{k=1}^{d}y_{k}=0,\text{ \ }\left\|y_{k}\right\|=1,\text{ \ }y_{k_{1}}\cdot y_{k_{2}}=-\frac{1}{d-1},\text{ \ }\forall k_{1}\neq k_{2},$ (A5)
let us introduce the set Ωy~={y~k=Uykd21,
k=1,..,d}
[…]
of vectors $\widetilde{y}_{k}$ defined to each $y_{k}$ due to the action of the unitary operator $\mathrm{U}:=\sum_{n}|e_{n}\rangle\langle g_{n}|,$ where $e_{n}$ and $g_{n}$ are eigenvectors in (A3). Vectors $\widetilde{y}_{k}$ satisfy relations (A5) and, therefore, projection
$\Pi_{\Omega_{\widetilde{y}}}=\frac{d-1}{d}\sum_{k=1}^{d}|\widetilde{y}_{k}\rangle\langle\widetilde{y}_{k}|=\mathrm{U}\Pi_{\Omega_{y}}\mathrm{U}^{\dagger}=\sum_{n=1}^{d-1}|e_{n}\rangle\langle e_{n}|$ (A6)
is included into the set of projections over which the maximum in (46) is considered. We have
$\mathrm{tr}[A\Pi_{\Omega_{\widetilde{y}}}]$ $=\mathrm{tr}\left[\sum_{n=1}^{N}\zeta_{n}|e_{n}\rangle\langle e_{n}|\left(\mathrm{U}\Pi_{\Omega_{y}}\mathrm{U}^{\dagger}\right)\right]$ (A7)
$=\mathrm{tr}\left[\sum_{n=1}^{N}\zeta_{n}|e_{n}\rangle\langle e_{n}|\text{ }\sum_{n^{\prime}=1}^{d-1}|e_{n^{\prime}}\rangle\langle e_{n^{\prime}}|\right]=\sum_{n=1}^{d-1}\zeta_{n}.$
Eqs. (A4), (A7) prove relation (46).
6.1 Examples
We proceed to present the explicit examples of vectors in a set $\Omega_{\widetilde{y}}$ on which the maximum (46) is attained. Consider first the case of an even $d\geq 2$. Introduce vectors
 $\text{\ }\widetilde{y}_{k}$ $=\frac{1}{\sqrt{d-1}}\sum_{n=1}^{d-1}\alpha_{kn}e_{n},\text{ \ \ }k=1,...,d,\text{ \ \ }\sum_{k=1}^{d}\alpha_{kn}=0,\text{ \ \ }n=1,...,d-1,$ (A8)
$\alpha_{1n}$ $=1,\text{ \ \ }\alpha_{kn}=\pm 1,\text{ \ \ }k=2,...,d,$
with the following properties
$\sum_{n=1}^{d-1}\alpha_{kn}\alpha_{k^{\prime}n}$ $=-1,\text{ \ \ }\forall k\neq k^{\prime}\in\left\{1,...,d\right\},$ (A9)
$\sum_{k=1}^{d}\alpha_{kn}\alpha_{kn^{\prime}}$ $=0,\text{ \ \ }\forall n\neq n^{\prime}\in\{1,...,d-1\}.$
Here, $e_{n}$ are eigenvectors of a positive operator $A.$ By (A8) and (A9) these vectors satisfy relations (A5), namely:
$\sum_{k=1}^{d}\widetilde{y}_{k}$ $=\frac{1}{\sqrt{d-1}}\sum_{n=1}^{d-1}\left(\sum_{k=1}^{d}\alpha_{kn}\right)e_{n}=0,$ (A10)
$\widetilde{y}_{k}\cdot\widetilde{y}_{k^{\prime}}$ $=\frac{1}{d-1}\sum_{n,n^{\prime}=1}^{d-1}\alpha_{kn}\alpha_{k^{\prime}n^{\prime}}\left(e_{n}\cdot e_{n^{\prime}}\right)$
$=\frac{1}{d-1}\sum_{n=1}^{d-1}\alpha_{kn}\alpha_{k^{\prime}n}=-\frac{1}{d-1},\text{ \ }\forall k\neq k^{\prime}.$
Moreover,
$\Pi_{\Omega_{\widetilde{y}}}=$ $\frac{d-1}{d}\sum_{k=1}^{d}|\widetilde{y}_{k}\rangle\langle\widetilde{y}_{k}|\text{ \ }=\frac{1}{d}\sum_{k=1}^{d}\left(\sum_{n,n^{\prime}\text{ }=1,...,(d-1)}\alpha_{kn}\alpha_{kn^{\prime}}|e_{n}\rangle\langle e_{n^{\prime}}|\right)$ (A11)
$=\sum_{n=1}^{d-1}|e_{n}\rangle\langle e_{n}|\text{ }+\text{ }\frac{1}{d}\sum_{n,n^{\prime}=1}^{d-1}\left(\sum_{k=1}^{d}\alpha_{kn}\alpha_{kn^{\prime}}\right)|e_{n}\rangle\langle e_{n^{\prime}}|$
$=\sum_{n=1}^{d-1}|e_{n}\rangle\langle e_{n}|.$
Since all $\alpha_{1n}=1,$ $\forall n=1,...,d-1,$ from relation $\sum_{k=1}^{d}\alpha_{kn}=0$ in (A8) it follows, for a fixed even $d$, the number $Z_{d}(+)$ of positive $\alpha_{kn}=+1$ in each $\widetilde{y}_{k}$, $k\neq 1,$ in (A8) is equal to $Z_{d}(+)=\frac{d-2}{2}$ whereas the number $Z_{d}(-)$ of negative $\alpha_{kn}=-1,k\neq 1,$ is equal to $\frac{d}{2}.$
For an odd $d>2$, we take the unit vectors
$y_{1}^{\prime}$ $=e_{1},\text{ \ \ \ }y_{2}^{\prime}=-\frac{1}{d-1}e_{1}+\frac{\sqrt{d}}{d-1}\sum_{n=2}^{d-1}e_{n},$ (A12)
$y_{k}^{\prime}$ $=-\frac{1}{d-1}e_{1}+\frac{\sqrt{d}}{d-1}\sum_{n=2}^{d-1}\alpha_{kn}e_{n},\text{ \ \ }\alpha_{kn}=\pm 1,\text{ \ }k=3,...,d,$
$\sum_{k=2}^{d}\alpha_{kn}$ $=0,\text{ \ \ }n=2,...,(d-1),\text{ \ \ }\sum_{d=2}^{d-1}\alpha_{k_{1}n}\alpha_{k_{2}n}=-1,\text{ \ \ }k_{1}\neq k_{2}\in\{2,...,d\},$
$\sum_{k=2}^{d}\alpha_{kn_{1}}\alpha_{kn_{2}}$ $=0,\text{ \ \ }\forall n_{1}\neq n_{2}\in\{2,...,d-1\}.$
In each vector $y_{k\geq 3}^{\prime},$ the number $Z_{d}(+)$ of positive $\alpha_{kn}=+1$ is equal to $Z_{d}(+)=\frac{d-3}{2}$ and the number $Z_{d}(-)$ of $\alpha_{kn}=-1,k\neq 1,$ is equal to $\frac{d-1}{2}.$ For these vectors,
$\sum_{k=1}^{d}y_{k}^{\prime}=0,\text{ \ \ \ }y_{k_{1}}^{\prime}\cdot y_{k_{2}}^{\prime}=-\frac{1}{d-1},$ (A13)
and
$\sum_{k}|y_{k}^{\prime}\rangle\langle y_{k}^{\prime}|\text{ }$ $=|e_{1}\rangle\langle e_{1}|\text{ }+\text{ }\frac{d-1}{(d-1)^{2}}|e_{1}\rangle\langle e_{1}|$ (A14)
$-\frac{\sqrt{d}}{(d-1)^{2}}\sum_{n=2}^{d-1}\left(\sum_{k=2}^{d}\alpha_{km}\right)\left\{|e_{1}\rangle\langle e_{n}|\text{ }+|e_{n}\rangle\langle e_{1}|\right\}$
$+\frac{d}{(d-1)^{2}}\sum_{n_{j}=2}^{d-1}\left(\sum_{k=2}^{d}\alpha_{kn_{1}}\alpha_{kn_{2}}\right)|e_{n_{1}}\rangle\langle e_{n_{2}}|$
$=\frac{d}{d-1}|e_{1}\rangle\langle e_{1}|\text{ }+\text{ }\frac{d}{(d-1)}\sum_{n=2}^{d-1}\text{ }|e_{n}\rangle\langle e_{n}|,$
so that
$\Pi_{\Omega_{{y^{\prime}}}}=\frac{d-1}{d}\sum_{k=1}^{d}|y_{k}^{\prime}\rangle\langle y_{k}^{\prime}|=\sum_{n=1}^{d-1}\text{ }|e_{n}\rangle\langle e_{n}|.$ (A15)
Let us now present the explicit examples of vectors, satisfying relations (A8) and (A12).
  1. 1.
    $d=2:$  in this case, $Z_{2}(+)=0,$ $Z_{2}(-)=1,$ and the vectors $\widetilde{y}_{1}=e_{1},$ $\widetilde{y}_{2}=-e_{1}$, satisfy all the relations in(A8).
  2. 2.
    $d=4:$ in this case $Z_{4}(+)=1,Z_{4}(-)=2$ and the vectors
    $\widetilde{y}_{1}$ $=\frac{1}{\sqrt{3}}(e_{1}+e_{2}+e_{3}),\text{ \ \ }\widetilde{y}_{2}=\frac{1}{\sqrt{3}}(e_{1}-e_{2}-e_{3}),$ (A16)
    $\widetilde{y}_{3}$ $=\frac{1}{\sqrt{3}}(-e_{1}+e_{2}-e_{3}),\text{\ \ \ }\widetilde{y}_{4}=\frac{1}{\sqrt{3}}(-e_{1}-e_{2}+e_{3}),$
    satisfy all relations in (A8).
  3. 3.
    $d=6:$ in this case, $Z_{6}(+)=2,$ $Z_{6}(-)=3$, and the vectors
    $\widetilde{y}_{1}$ $=\frac{1}{\sqrt{5}}(e_{1}+e_{2}+e_{3}+e_{4}+e_{5}),\text{ \ }\widetilde{y}_{2}=\frac{1}{\sqrt{5}}(e_{1}+e_{2}-e_{3}-e_{4}-e_{5}),\text{\ }$ (A17)
    $\widetilde{y}_{3}$ $=\frac{1}{\sqrt{5}}(-e_{1}-e_{2}-e_{3}+e_{4}+e_{5}),\text{\ \ \ }\widetilde{y}_{4}=\frac{1}{\sqrt{5}}(-e_{1}+e_{2}+e_{3}-e_{4}-e_{5}),$
    $\widetilde{y}_{5}$ $=\frac{1}{\sqrt{5}}(-e_{1}-e_{2}+e_{3}+e_{4}-e_{5}),\text{\ \ \ }\widetilde{y}_{6}=\frac{1}{\sqrt{5}}(e_{1}-e_{2}-e_{3}-e_{4}+e_{5}),$
    satisfy relations in (A8).
  4. 4.
    $d=3:$ then the vectors
    $y_{1}^{\prime}=e_{1},\text{ \ \ \ }y_{2}^{\prime}=-\frac{1}{2}e_{1}+\frac{\sqrt{3}}{2}e_{2},\text{ \ \ }y_{3}^{\prime}=-\frac{1}{2}e_{1}-\frac{\sqrt{3}}{2}e_{2}\text{,}$ (A18)
    satisfy relations (A12).
  5. 5.
    $d=5$: then in each $y_{k\geq 3}^{\prime}$, the number $Z_{5}(+)$ of $\alpha_{kn}=1$ is equal to $1$ and the number $Z_{5}(-)_{-}$ of $\alpha_{kn}=-1$ is equal to $2$ The vectors
    $y_{1}^{\prime}$ $=e_{1},\ \ \ y_{2}^{\prime}=-\frac{1}{4}e_{1}+\frac{\sqrt{5}}{4}(e_{2}+e_{3}+e_{4}),\ \ y_{3}^{\prime}=-\frac{1}{4}e_{1}-\frac{\sqrt{5}}{4}\left(e_{2}-e_{3}-e_{4}\right),$ (A19)
    $y_{4}^{\prime}$ $=-\frac{1}{4}e_{1}-\frac{\sqrt{5}}{4}\left(-e_{2}-e_{3}+e_{4}\right),\text{ \ \ \ \ }y_{5}^{\prime}=-\frac{1}{4}e_{1}-\frac{\sqrt{5}}{4}\left(-e_{2}+e_{3}-e_{4}\right),$
    satisfy the relations in (A12).